Jan mixes 1 1/3 cups of white flour, 1 3/4 cup of whole wheat flour, and 2/3 cup of rye flour. About what fraction of the mixture is whole wheat flour?
Answer by richwmiller(17219)

4/3+7/4+2/3=
(16+21+8)/12
7/4=21/12 whole wheat
45/12=total
21/45=whole wheat /total
7/15 is whole wheat